# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v2

@@ 4.2 Main Results
Like a miracle: LLM enables English CLIP to learn Chinese without exposure to Chinese data.
To validate the knowledge transfer capabilities of the LLM, we designed an out-of-distribution (OOD) experiment by conducting an image-text retrieval task in a completely unfamiliar language. This is an especially challenging experiment because all of our training was performed on purely English text, yet now we are testing the model on Chinese data. As shown in Table 3, models like EVA02 CLIP and Jina CLIP, which performed well on English datasets, achieved near-zero accuracy on this task. However, the magical power of LLM2CLIP became evident: the LLM’s capabilities allowed the model to achieve impressive performance on Chinese retrieval tasks, even surpassing models that were originally trained on hundreds millions of Chinese data, such as Wukong (Gu et al., 2022) and CN-CLIP (Yang et al., 2022). This result once again demonstrates that LLM2CLIP can effectively integrate the inherent abilities of the LLM into CLIP, enabling it to handle tasks far beyond its original scope.
LLM2CLIP Can Improve the Performance of VLLM.
As shown in Table 7, we conducted experiments using the Llava 1.5 Liu et al. (2023) VLLM (Vision-Language Large Model) training framework. Llava incorporates a CLIP visual encoder into the LLM for multimodal instruction learning, meaning the quality of the visual encoder can significantly impact Llava’s performance. We compared the original CLIP with a version enhanced by our LLM2CLIP fine-tuning, running two versions of the experiment according to Llava’s official implementation for a fair comparison. The results showed that in over 87.5% of the benchmarks, we achieved substantial performance improvements, with the remaining benchmarks showing results that were very close. This demonstrates the potential of LLM2CLIP for complex image reasoning and related tasks.
@@ 4.3 Ablation
With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions. In  Table 5, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance. The “Ratio” column represents the proportion of dense captions used during training. Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better. For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results. Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst. It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal. Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data, striking a balance between both short-text and long-text retrieval tasks.
@@ 5 Limitations and Future Work
While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential. LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains. However, we did not intentionally select fine-tuning data based on specific characteristics. By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks. Moreover, in this work, we froze the gradients of the LLM during fine-tuning to maintain a large batch size for CLIP training. However, it would be worth exploring approaches that allow for updating the LLM’s gradients, as there may be better engineering practices that can balance the two objectives. Re-training LLM2CLIP from scratch on datasets like Laion-2B (Schuhmann et al., 2022) and Recaption-1B (Li et al., 2024b) is also a promising direction that we did not pursue due to time constraints.
@@ 6 Conclusion
This paper presents a method that enables LLMs to assist in CLIP training. Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications. The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training. Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models. We experimented with LLM2CLIP on both CLIP and EVA models and validated it across different Llama3 models, including 1B, 8B, and 13B, all of which yielded transformative improvements in CLIP performance. LLM2CLIP offers several advantages, such as seamless compatibility with long and complex text inputs, and the ability to incorporate the open-world knowledge of LLMs into CLIP training — as demonstrated by its ability to match Chinese text even when trained solely on English data. We firmly believe that the potential of LLMs extends far beyond what has been shown, and we hope this work serves as a starting point for using LLMs to enhance various vision foundation models, ultimately spreading their general capabilities and open-world knowledge across a wide range of modalities and applications.

### v3

@@ 4.2.2 Cross-Lingual Retrieval
LLM Enables English CLIP to Learn Chinese Without Exposure to Chinese Data.
To test LLM knowledge transfer, we conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. As shown in Table 3, EVA02 CLIP and JinaCLIP showed near-zero performance, while LLM2CLIP achieved remarkable results, even outperforming models trained on extensive Chinese data like Wukong [12] and CN-CLIP [42]. This demonstrates the power of LLMs in transferring knowledge to cross-modal spaces.
4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset.
We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe.
Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
4.2.4 Multimodal Large Language Models Performance
Evaluation Dataset.
Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
LLM2CLIP Improves VLLM Performance.
As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Main Conclusions
Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
Directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance, as shown by the red-highlighted numbers in Table 7 and Table 6, with average performance drops of 10%-20%. Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features lack sufficient discriminability for captions. These results clearly demonstrate that such an approach imposes a significant burden on CLIP’s learning process.
To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning. Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora(same as LLM2Vec) and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss. As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks. This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features. These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
LLM’s Built-in Knowledge Directly Benefits CLIP.
In Table 7, we observe that directly using LLM2CLIP leads to significant performance improvements even with short caption datasets, which is quite challenging, as CLIP has already seen most of this training data. However, the results from LLM2CLIP with +Llama3-8B-CC still show a clear performance gain, which can be attributed purely to the improvement in the text encoder. On the other hand, +Jina-Bert fails to achieve similar results, and in fact, leads to a performance decline.
LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications. Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT. To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions. As shown in Table 7, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
Larger LLMs Lead to Better Performance.
LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training. We thoroughly investigate the performance differences that various versions of LLMs bring to CLIP. In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact. We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder. As expected, within the same Llama family, the 8B model significantly outperfoWrms the 1B model. Mistral Memo also shows a slight improvement over the 8B model. It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP.
@@ 4.4 Ablation
With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions. In  Table 8, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance. The “Ratio” column represents the proportion of dense captions used during training. Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better. For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results. Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst. It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal. Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data.
@@
Adaptor design.
In Table 9, we explore the impact of the Adaptor and its layer count. We find that adding at least one layer of the Adaptor is beneficial, resulting in an average performance improvement of two points. However, increasing it to four layers provides minimal additional gains.
Caption Contrastive Fine-tuning Ablation.
In Table 10, we explore the impact of the key steps in LLM2Vec’s Caption Contrastive fine-tuning, namely MNTP and bidirectional attention. Through comparisons, we observe that MNTP offers almost no positive effect, while bidirectional attention provides some performance gains but is not decisive. Therefore, we conclude that the most significant contribution comes from the supervised SimCSE contrastive learning applied to the captions.
@@ 5 Limitations and Future Work
This paper focuses solely on the impact of LLMs on CLIP. In reality, there are many other areas where text can play a transformative role, such as diffusion models [34, 4, 39], Segment Anything [16] and GroundingDINO [25], both of which could potentially see significant performance improvements by applying our method.
While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential. LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains. However, we did not intentionally select fine-tuning data based on specific characteristics. By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks.
@@ 6 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications. The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training. Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models.